A mixture of three waves lambda1,2,3 are made incident on metal surface of threshold wavelength lambda 2.if no. of photons of each wavelength is 10^6 and efficiency of photolectric effect is 1%.calculate no. of photoelectrons emitted?

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Efficient of photoelectric effect is the ratio of number of photoelectrons ejected from the metal surface and the number of incident photons,

Hence the number of photoelectrons emitted in each case, regardless of the wavelength of the incident waves is0.01\times 10^6=10^4